### v1

# Paint Outside the Box: Synthesizing and Selecting Training Data for Visual Grounding

Abstract
Visual grounding aims to localize the image regions based on a textual query. Given the difficulty of large-scale data curation, we investigate how to effectively learn visual grounding under data-scarce settings in this paper. To address data scarcity, we propose a novel framework, POBF (Paint Outside the Box, then Filter). POBF synthesizes images by inpainting outside the box, tackling a label misalignment issue encountered in previous works. Furthermore, POBF leverages an innovative filtering scheme to identify the most effective training data. This scheme combines a hardness score and an overfitting score, balanced by a penalty term. Experimental results show that POBF achieves superior performance across four datasets, delivering an average improvement of 5.83% and outperforming leading baselines by 2.29% to 3.85% in accuracy. Additionally, we validate the robustness and generalizability of POBF across various generative models, data ratios, and model architectures.
1 Introduction
Visual grounding [39, 41, 93, 11, 29, 90, 63, 12] is a fundamental task in vision-language understanding. The objective is to locate a specific image region based on a free-form textual query. However, constructing visual grounding datasets [92, 47, 50, 30] with dense region-text annotations is both time-intensive and costly, which hinders the progress in this field. Motivated by the difficulty of large-scale data curation, we investigate the feasibility of effectively learning visual grounding under data-scarce settings, such as with access to only 1% training data.
With recent advances in image generation and captioning models [36, 54, 55, 40], a natural approach to address data scarcity is to generate training data. The advanced generative models, extensively pretrained on large-scale real-world data, have captured the correspondence between natural images and their textual descriptions. It is probable to benefit visual grounding by distilling the knowledge from generative models through training data generation. Recent works have explored the generation of densely annotated data to train models for fine-grained understanding tasks, such as object detection and image segmentation [96, 14, 83, 16, 38, 15, 7, 4]. However, these methods rely on specialized networks trained on extensive dense annotations, which hinders their flexibility and generalizability.
To address this limitation, we propose leveraging off-the-shelf generative models, trained solely on image-caption pairs in a weakly-supervised manner, to synthesize training data. Our method offers two advantages. First, the use of the off-the-shelf generative models allows our method to generalize across a wider range of domains, since the generative models are trained on web-scale image-caption data. Second, we can benefit from the rapid progress in generative models, incorporating the latest advancements into our data generation pipeline without additional training on dense annotations.
Figure 1: Examples of generated images from different methods. X-Paste [96] synthesizes unnatural images. Gen2Det [64] exhibits the label misalignment issue. GeoDiffusion [7] either fails to generate the correct object or introduces excessive artifacts. In contrast, our method synthesizes high-quality images with accurately aligned bounding boxes.
Without being trained on dense annotations, the off-the-shelf generative models struggle to directly generate training data with dense annotations. To address this, we utilize the region-text annotations from limited real data to generate training data through inpainting. However, naively inpainting the object within the bounding box [64] can lead to label misalignment issues. Specifically, the generated object may not spatially align well with the box or even fail to appear entirely. Figure 1 shows the examples. The synthetic ‘right cow’ in 1(a) misaligns with the box, and the ‘batter’ in 1(b) disappears. To mitigate this problem, we propose a strategy that preserves the content within the box while generating the background outside. Figure 1 shows we can generate high-quality synthetic images that align well with the original annotations.
Moreover, training neural networks with synthetic data presents another challenge: not all samples contribute positively to model performance [4, 10, 46, 45]. To identify the most effective training data, we propose a filtering scheme consisting of three components: a hardness score, an overfitting score, and a penalty term. The hardness score prioritizes the easiest samples, as it is challenging to model outliers when the basics are not adequately captured in our data-scarce regime [62]. The overfitting score removes samples exacerbating overfitting outside the box, as gigantic models are prone to overfitting on unintended features (known as shortcut learning[18]), especially with limited training data. Furthermore, to manage the trade-off between two scores, we introduce a penalty term. Finally, we combine the selected synthetic data with real data as the training data to train the model.
We propose a novel framework named POBF (Paint Outside the Box, then Filter). This approach leverages generative models to inpaint and caption images, thereby expanding the limited training set effectively. Additionally, it incorporates a filtering scheme to select the most effective synthetic data for training. Our contributions are as follows:
  • We address a challenging data-scarce scenario in visual grounding by proposing POBF, an innovative framework that synthesizes images by inpainting outside the box. POBF uses off-the-shelf data generators pretrained solely on image-caption pairs, which makes it more flexible and adaptable.
  • We propose a novel filtering scheme tailored to our data-scarce setting to identify the most effective synthetic training samples. This scheme comprises a hardness score, an overfitting score, and a penalty term.
  • Experimental results demonstrate that our proposed framework yields a 5.83% improvement across four datasets. Additionally, POBF surpasses leading baselines, such as X-Paste [96] and GeoDiffusion [7], with an accuracy increase ranging from 2.29% to 3.85%. We validate the robustness and generalizability of POBF across generative models, data ratios, and model architectures.
2 Problem Definition
Let $I$ and $T$ denote an input image and a natural language description, respectively. Visual grounding is the task of identifying the target region $R\subseteq I$ within the image that the description refer to. Typically, $R$ is represented as a bounding box $B=(x,y,w,h)$, where $(x,y)$ specifies the box center coordinates, and $w$ and $h$ define its width and height. In this paper, we formulate the problem of learning visual grounding from data-scarce scenarios. Given a dataset $D$, we aim to train a model effectively using only a small subset comprising $N$of the entire data, where $N\%\ll 1$.
3 Methodology
Figure 2: Overview of our proposed framework POBF, which consists of four key steps: data generation, teacher training, data filtering, and student training. Here, $(I,T,B)$ represents a real data sample, while $T^{\prime}$ and $I^{\prime}_{j}$ denote the generated caption and image, respectively. $\mathcal{G}_{T}$ and $\mathcal{G}_{I}$ are the generative models. $\mathcal{T}$ and $\mathcal{S}$ refer to the teacher and student model, repectively.
In this paper, we introduce a novel, model-agnostic framework called POBF (Paint Outside the Box, then Filter). An overview of our pipeline is provided in Figure 2. Our approach begins with a data generation process, where we synthesize images and captions using image inpainting and captioning techniques. This process utilizes the image generator and captioner pretrained on large-scale image-caption pairs. Next, we train a teacher model using the limited real data. We then apply a filtering scheme to select effective synthetic images. Specifically, the well-trained teacher network assigns a hardness score, an overfitting score, and a penalty term to each synthetic image. Finally, we retain the synthetic images with high scores and combine them with real data to train the student model. The proposed POBF framework is simple, effective, and does not rely on any auxiliary models pretrained with dense annotated data.
3.1 Cross-modality Generation
Image Generation with Inpainting. We begin by synthesizing images through image inpainting, a process that reconstructs masked regions within an image based on a given prompt, while preserving the rest of the image. To address the label misalignment problem, we propose inpainting outside of the bounding box instead. For each data point ($I,T,B$), we first use a captioning model $\mathcal{G}_{T}$ to generate a caption $C$ for $I$. The bounding box $B$, image $I$, and caption $C$ are then input into the inpainting model $\mathcal{G}_{I}$, yielding an generated image $I^{\prime}$. As shown in Figure 2, the red bus within the bounding box remains unchanged, allowing it to continue serving as the ground truth. The synthetic image $I^{\prime}$ can be paired with the original text $T$ and bounding box $B$ to form a new training sample ($I^{\prime},T,B$). By generating $K$ synthetic images per real sample, denoted as $\{I^{\prime}_{1},I^{\prime}_{2},\dots,I^{\prime}_{K}\}$, the training dataset effectively increases in size by a factor of $K$.
We utilize Stable Diffusion XL [54] as the inpainting model, which is pretrained on image-caption pairs with the objective of predicting masked regions. To increase diversity, we set a large strength parameter, which controls the amount of noise added to the base image thereby affects the similarity of the output to the base image.
Text Generation with Captioning. To augment the original text $T$, we feed the cropped image within the bounding box $B$ to the captioning model $\mathcal{G}_{\text{T}}$, which generates a new caption $T^{\prime}$ describing the object. The generated caption $T^{\prime}$ can be paired with $B$ and either the original image $I$ or the generated image $I^{\prime}$ to create a new training sample. In this work, we employ BLIP [36] as the image captioner, with nucleus sampling [24] to improve diversity.
3.2 Filtering Scheme
Recognizing that not all synthetic samples contribute positively to model performance, we propose a novel filtering scheme to select the most effective images for training. Initially, we use the accessible real training dataset to train a teacher model $\mathcal{T}$. Then, $\mathcal{T}$ evaluates synthetic images based on two scores and one penalty term to determine their effectiveness as training samples. We provide an illustration in Figure 3 to demonstrate the input used when calculating each score for a generated image. For clarity in the following discussions, we assume a real data sample $(I,T,B)$ and denote the set of images generated from it as $\{I^{\prime}_{1},I^{\prime}_{2},\dots,I^{\prime}_{K}\}$.
Hardness Score. We propose to filter harmful synthetic images via computing hardness. Intuitively, in our data-scarce regime, it is difficult to model outliers effectively as fundamental patterns may not be fully captured [62]. Since the model require easiest examples to provide coarse-grained information about the target function, we prioritizes synthetic images with lower losses. We define the hardness score for each of the synthetic image $I_{j}^{\prime}$, where $j\in\{1,2,\dots,K\}$, as follows:
$S_{1}(I_{j}^{\prime})=\text{IOU}(\mathcal{T}(I_{j}^{\prime},T),B)$ (1)
where $\mathcal{T}(I_{j}^{\prime},T)$ is the predicted bounding box, and IOU (Intersection over Union) is the ratio of the intersection of two areas to their combined area.
Overfitting Score. Additionally, we introduce a novel score to filter out synthetic images that may hinder training by exacerbating overfitting. Given the gigantic model architecture relative to the limited number of real data, the model is prone to overfit by relying on image background features rather than the actual object in the bounding box. To mitigate this issue, we assess the potential of each generated image to contribute to overfitting. Specifically, we modify each generated image by masking the area within the bounding box and input this variant into $\mathcal{T}$. If $\mathcal{T}$ can still correctly predict the bounding box, it indicates that $\mathcal{T}$ may leverage unintended features present in the generated background. Incorporating such data into training dataset risks the model learning these undesirable features, thereby exacerbating overfitting. To prevent this, we define an overfitting score to filter these problematic samples. The score for each generate image $I_{j}^{\prime}$, where $j\in\{1,2,\dots,K\}$, is defined as:
$S_{2}(I_{j}^{\prime})=1-\text{IOU}(\mathcal{T}(\text{Mask}(I_{j}^{\prime},B),T),B)$ (2)
where $\text{Mask}(I_{j}^{\prime},B)$ is a operation that sets all pixel values within the bounding box $B_{i}$ to zero.
Figure 3: An illustration of the process of calculation each score for a generated image.
Penalty Term. To balance between preserving easy samples and the mitigation of overfitting risk, we introduce a penalty term. The rationale is that overly easy images may not challenge the model enough to learn complex patterns. Conversely, images where the background outside the bounding box does not offer any useful information may introduce model bias. To balance between the first two scores, we define a penalty term for each generate image $I_{j}^{\prime}$, where $j\in\{1,2,\dots,K\}$, as follows:
$P(I_{j}^{\prime})=\text{IOU}(\mathcal{T}(I_{j}^{\prime},\emptyset),B)$ (3)
where $\emptyset$ denote an empty textual string.
The final score for a generated image $I_{j}^{\prime}$ is calculated as:
$\text{Score}(I_{j}^{\prime})=\lambda_{1}S_{1}(I_{j}^{\prime})+\lambda_{2}S_{2}(I_{j}^{\prime})+\lambda_{P}P(I_{j}^{\prime})$ (4)
where $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{P}$ are hyperparameters controlling the weights of each component.
Out of the $K$ generated images $\{I^{\prime}_{1},I^{\prime}_{2},\dots,I^{\prime}_{K}\}$ from each real data point ($I,T,B$), we retain one synthetic image $I^{\prime}_{m}$ with the highest score, where $m=\arg\max_{j}\text{Score}(I^{\prime}_{j})$ and $j\in\{1,2,\dots,K\}$. This yields a synthetic data point ($I^{\prime},T,B$). By incorporating the selected synthetic data points into the training set, we effectively double its size. Finally, we then train a student model $\mathcal{S}$. During training, the real text $T$ is randomly replaced with the corresponding generated caption $T^{\prime}$ with probability $q$.
4 Experiments
Dataset. RefCOCO [92], RefCOCO+ [92], and RefCOCOg [47, 50] datasets are derived from MSCOCO 2014. Following the common splits used in previous studies [11, 12], we report performance on the testA (containing multiple people) and the testB (containing multiple instances of other objects) for RefCOCO and RefCOCO+, and the umd-test splits for RefCOCOg [50]. The ReferItGame dataset [30] is collected through an online two-player game. We report performance on the testset of ReferItGame. The statistics for these four datasets are presented in Table 1. For each dataset, we randomly sample a subset comprising 1% of the entire dataset to simulate data-scarce scenarios.
Table 1: Statistics of visual grounding datasets.
Dataset Images Texts Objects
RefCOCO [92] 19,994 142,210 50,000
RefCOCO+ [92] 19,992 141,564 49,856
RefCOCOg [47, 50] 25,799 95,010 49,822
ReferItGame [30] 19,894 130,525 96,654
Table 2: Performance comparison with state-of-the-art baselines. All methods employ the proposed filtering scheme.
Method Dataset RefCOCO RefCOCO+ RefCOCOg ReferIt Avg
Real Syn TestA TestB TestA TestB Test Test
Results on One Sampled Dataset
Real 44.26 34.34 23.58 17.70 24.87 21.87 27.78
GeoDiffusion [7] 47.22 34.90 28.01 18.83 25.39 24.20 29.76+1.98
Gen2Det [64] 46.61 35.19 28.85 20.65 26.10 25.00 30.40+2.62
X-Paste [96] 47.85 39.01 27.59 21.25 27.08 25.16 31.32+3.54
Ours 52.77 40.96 31.37 21.25 29.98 25.28 33.61+5.83
Average Results across Three Sampled Datasets
Real 43.64$\pm$0.98 34.89$\pm$0.50 25.15$\pm$1.43 18.02$\pm$1.48 24.92$\pm$1.05 22.94$\pm$0.97 28.26
Ours 51.25$\pm$1.49 39.62$\pm$1.16 32.15$\pm$0.68 22.38$\pm$1.01 30.73$\pm$0.69 25.10$\pm$0.30 33.54+5.28
Baselines. We compare our method with three recent data generation techniques [96, 7, 16] designed for fine-grained understanding tasks: (i) X-Paste [96] uses an ensemble of pretrained segmentation models to identify instances from generated images, which are then copied into real images. (ii) Gen2Det [64] employs a grounded editing diffusion model to edit objects in real images. (iii) GeoDiffusion [7] utilizes geometric control via text prompts to enhance text-to-image diffusion models for object detection data generation. For a fair comparison, we refrain from using the models that have been pretrained on a large scale with fine-grained annotations. Implementation details can be found in the supplementary material.
Evaluation Metric. Following previous works [11, 12], we use top-1 accuracy as the evaluation metric, where a prediction is considered correct if the IoU between the predicted bounding box and the ground-truth exceeds 0.5.
Implementation Details. For both the teacher and student networks, we utilize TransVG [11], one of the most popular models for visual grounding, known for being the first transformer-based architecture designed without any bells or whistles. We utilize the Stable Diffusion XL model [54] for image inpainting, with 45 denoising steps, a strength of 0.9, and a guidance scale of 7.5. BLIP ViT-L [36] is employed as the image captioner, with nucleus sampling ($p$=0.9).The hyperparameters $\lambda$ in filtering scheme is tuned using grid search on the validation set, with values of 1.0 or 0.5. For each real data point, we generate 4 synthetic images and apply our proposed selection strategy to retain one with highest score. To ensure a fair comparision, we adopt a consistent training schedule (batch size × steps = 64 × 6500) across all experiments, following previous works [95, 73, 83].
4.1 Main Results
Comparison with SOTA Baselines. Table 2 presents a comparative analysis of our proposed method against several state-of-the-art baselines across multiple benchmark datasets. We have following observations. (1) Our method outperforms the baseline trained solely on real data by an average of 5.83%, highlighting the effectiveness of our inpainting strategy and filtering scheme. (2) Models utilizing synthetic data generated by various baselines show at least a 1.98% improvement compared to the model trained exclusively on real data, demonstrating the substantial value of synthetic data in enhancing performance. (3) Our proposed method consistently outperforms all other baselines, achieving the highest accuracy across all datasets. The best-performing baseline, X-Paste, achieves an average score of 31.32%, but our approach surpasses it by 2.29%, achieving the state-of-the-art performance. Notably, our model reaches an impressive 52.77% accuracy on RefCOCO TestA and 31.37% on RefCOCO+ TestA, significantly outperforming other baselines. Additionally, our method achieves gains of 2.90% on RefCOCOg Test and 0.18% ReferIt compared to X-Paste, demonstrating robustness and generalizability across diverse datasets. We hypothesize that the relatively modest performance gain on ReferIt is due to the prevalence of small objects in this dataset. The inpainting process for such large background areas introduces noticeable artifacts, which may adversely affect model performance. In conclusion, our approach sets a new benchmark for visual grounding in our data-scarce scenario, representing a significant advancement over existing SOTA models.
Table 3: Ablations on synthetic data and filtering scheme. $\text{Syn}_{I}$ and $\text{Syn}_{T}$ denote synthetic images and texts, respectively.
Training Data Filtering RefCOCO RefCOCO+ RefCOCOg ReferIt Avg
Real $\text{Syn}_{I}$ $\text{Syn}_{T}$ $S_{1}$ $S_{2}$ $P$ TestA TestB TestA TestB Test Test
the Effectiveness of the Synthetic Data
44.26 34.34 23.58 17.70 24.87 21.87 27.78
49.97 39.16 29.01 20.56 26.67 22.49 31.31+3.53
47.00 36.19 25.36 18.74 26.33 22.13 29.29+1.51
the Effectiveness of the Filtering Scheme
50.13 38.86 29.10 20.10 28.31 23.39 31.65+3.87
50.66 39.56 30.63 21.76 30.18 24.50 32.88+5.10
50.29 39.67 29.95 21.17 28.42 24.67 32.36+4.58
51.15 39.24 30.22 21.44 29.70 25.01 32.79+5.01
51.94 39.75 30.65 22.44 29.41 24.30 33.08+5.30
52.77 40.96 31.37 21.25 29.98 25.28 33.61+5.83
Average Results on Three Randomly Sampled Datasets. Additionally, we sampled three different subsets from each real dataset to further evaluate performance. The results, presented in Table 2, demonstrate that despite performance variance due to the relatively small size of the sampled datasets, our method consistently outperforms the model trained solely on real data across all subsets. These findings indicate that our model’s performance is robust and generalizes well to variations in the data.
4.2 Ablation Study
Ablation on Synthetic Data. We begin by evaluating the effectiveness of synthetic data by separately using generated images and texts, without any filtering scheme. We observe that using only real data results in relatively low performance, with the model achieving an average accuracy of 27.78%. Introducing synthetic images leads to a significant improvement, boosting performance by 3.53% on average, demonstrating the effectiveness of our proposed inpainting strategy. When synthetic texts are applied, the improvement is modest at 1.51%, demonstrating the importance of generated data for performance gains.
Ablation on Filtering Scheme. Next, we assess the effectiveness of different components of our filtering scheme. The individual components—$S_{1}$, $S_{2}$, and $P$—result in performance gains of 1.23%, 0.71%, and 1.14%, respectively, compared to the variant without any filtering. These results demonstrate the effectiveness of each component in selecting valuable training data. Notably, the variant without the penalty term achieves an accuracy of 33.08%, outperforming all configurations that include just one filtering component. Combining synthetic data with the full filtering scheme leads to a substantial improvement, with the highest recorded gain of 1.96%. This configuration achieves the highest accuracy across all datasets, with a remarkable average accuracy of 33.61%, underscoring the effectiveness of our proposed filtering scheme.
In conclusion, all of these results demonstrate that both synthetic data and filtering schemes play crucial roles in enhancing model performance, with the combination of these elements yielding the best results.
Table 4: Results obtained by isolating scores without generated text. Results with a gray background indicate performance lower than the RANDOM baseline.
Filter RefCOCO RefCOCO+ RefCOCOg
TestA TestB TestA TestB Test
RANDOM 48.52 38.84 27.47 20.27 27.34
ALL 48.45 38.68 28.40 20.29 26.90
$S_{1}$ 49.94 40.14 29.53 20.67 28.10
$S_{2}$ 49.48 39.83 27.68 20.35 29.47
$P$ 50.66 39.47 28.50 20.41 27.81
Negative Score
1-$S_{1}$ 48.17 38.90 25.98 19.25 26.40
1-$S_{2}$ 48.22 38.51 26.86 19.60 25.54
4.3 Filtering Scheme Analysis
In this section, we isolate the each component in the filtering scheme without the influence of generated text, offering insights into their impact on model performance. We create two variants of our method: (i) The RANDOM method randomly select images as training data. (ii) The ALL baseline uses all synthetic data indiscriminately for training. We have following observations from Table 4.
(1) Despite the ALL method having access to more data than the RANDOM baseline, it sometimes underperforms. For instance, in RefCOCOg, there is a performance drop of 0.44%. This observation suggests that simply increasing the amount of synthetic data does not necessarily lead to performance gains. It underscores the importance of filtering out low-quality data to ensure the effectiveness of the training data. (2) Models using any of the three filtering components outperform the two baselines. This demonstrates the critical importance of leveraging $S_{1}$ and $S_{2}$ scores to select high-quality training data, aligning with our observations in Table 3 when synthetic texts are incorporated. (3) It is worth noting that different filtering scores achieve the highest results on different datasets due to the distinct data distributions. For instance, $S_{1}$ achieves peak performance on RefCOCO+ TestA with an accuracy of 29.53%, while $S_{2}$ yields the best performance on RefCOCOg with an accuracy of 29.47%. This variability suggests that each filtering criterion captures unique aspects of the data distribution, which are beneficial for model performance in specific datasets.
Interestingly, when data is filtered using negative values of $S_{1}$ or $S_{2}$, model performance drops below the random baseline. For example, using a 1-$S_{1}$ score reduces accuracy to 48.17% on RefCOCO TestA and 25.98% on RefCOCO+ TestA, compared to the random baseline scores of 48.52% and 27.47%, respectively. Similarly, filtering with 1-$S_{2}$ results in consistent performance declines across datasets, such as a 0.61% drop on RefCOCO TestA and a 1.80% drop on RefCOCOg. The observed performance drop when using negative scores suggests that our filtering scheme is effective in identifying valuable training data, as the introduction of poorly selected data (as indicated by negative scores) negatively impacts the model’s performance.
4.4 Cross-modality Generative Models
We evaluate whether POBF can seamlessly integrate with other state-of-the-art pretrained generative models. By default, POBF utilizes Stable Diffusion XL [54] as the image generator and BLIP ViT-L [36] as the image captioner. In this experiment, we test two alternative image generators: Stable Diffusion 2 [55] and Dreamshaper [^1] https://huggingface.co/Lykon/dreamshaper-8-inpainting. For image captioning, we explore replacements such as LLAVA [40] and ViT-GPT2222https://huggingface.co/nlpconnect/vit-gpt2-image-captioning. Table 5 shows the results. We observe that while the default configuration of POBF delivers robust performance, it does not always yield the highest accuracy. Specifically, model combinations using Dreamshaper as the image generator achieve superior outcomes, reaching an average accuracy of 27.63%. Furthermore, these configurations consistently outperform the baseline without our framework. These observations highlight the generalizability and effectiveness of our proposed framework.
Table 5: Performance comparison across generative models.
RefCOCO+ RefCOCOg Avg
Method TestA TestB Test
Real 23.58 17.70 23.95 21.74
Ours 31.37 21.25 29.98 27.53+5.79
Replace the Image Captioner with $\dots$
LLAVA 31.09 21.54 28.57 27.07+5.33
ViT-GPT2 29.95 21.22 29.90 27.02+5.28
Replace the Image Generator with $\dots$
SD 2 31.16 21.19 29.30 27.22+5.48
Dreamshaper 30.20 22.39 30.32 27.63+5.89
4.5 Ratios of Training Data
Table 6 presents the performance of our proposed framework under varying amounts of training data. We experiment with two scenarios: using 0.5% and 2.0% of the original training data. The results demonstrate that our method consistently provides substantial performance gains over the baseline. Specifically, with just 0.5% of the training data, the baseline model achieves an average accuracy of 13.94%. In contrast, our approach raises the average accuracy to 17.74%, representing a notable improvement of 3.80%. This highlights the effectiveness of our method in enhancing performance even under extremely limited data conditions. In the case of 2.0% training data, our method boosts the average accuracy by 3.99%. These consistent gains across both scenarios underscore the robustness of our approach, showcasing its potential for practical applications in data-scarce environments.
Table 6: Performance comparison across ratios of training data.
RefCOCO+ RefCOCOg Avg
Method TestA TestB Test
0.5% Training Data
Real 16.32 10.00 15.49 13.94
Ours 20.37 13.74 19.12 17.74+3.80
2.0% Training Data
Real 32.97 25.11 33.77 30.62
Ours 36.54 28.84 38.47 34.61+3.99
4.6 Model Architectures
We evaluate the proposed POBF framework across different model architectures to assess its generalizability and effectiveness. Specifically, we experiment with TransVG [11], Dynamic-MDTRE [60], and QRNet [90], each of which differs in terms of the number of parameters and architectural complexity. Table 7 presents the results. Although model performance generally increases with model size, our POBF framework consistently delivers extensive performance gains across all tested architectures. Specifically, Dynamic-MDTRE exhibits an more pronounced improvement, with average accuracy rising from 23.81% to 30.84%, resulting in a gain of 7.03% points. QRNet, the most complex model with 273M parameters, also benefits from a 4.41% point increase in average accuracy. The results highlight that our method is not tailored to a specific architecture but instead offers consistent performance benefits across a diverse set of models.
Table 7: Performance comparison across model architectures. #Paras denotes the number of model parameters.
RefCOCO+ RefCOCOg Avg
Method TestA TestB Test
TransVG [11] (#Paras = 149M)
Real 23.58 17.70 23.95 21.74
Ours 31.37 21.25 29.98 27.53+5.79
Dynamic-MDTRE [60] (#Paras = 160M)
Real 23.82 18.55 29.07 23.81
Ours 33.55 23.93 35.03 30.84+7.03
QRNet [90] (#Paras = 273M)
Real 32.32 23.82 30.30 28.81
Ours 37.29 26.65 35.73 33.22+4.41
5 Related Work
Visual Grounding. Early approaches addressed visual grounding in a two-stage framework [39, 93, 41], which heavily depended on the quality of proposals and large computational costs due to the large number of proposals. To overcome the limitations, more efficient one-stage methods have been developed [11, 29, 90, 63, 12, 76, 60, 77, 82]. TransVG [11] incorporates a DETR encoder for visual feature extraction and introduces the first transformer-based framework for visual grounding. QRNet [90], VG-LAW [63], and TransVG++ [12], integrate language-guided knowledge within the intermediate layers of visual backbones to compute query-dependent visual attention dynamically. As the first diffusion-based framework, LG-DVG [9] progressively reasons the queried object box by denoising a set of noisy boxes with the language guide. Recent works [21, 86] employ GradCAM [58] within the ALBEF model [35] to produce gradient-based visual explanations, represented as heatmaps for grounding outputs. Alternatively, other works [5, 6, 97] propose to train grounding multimodal large language models on extensive fine-grained labeled datasets. Previous works typically train the model on large-scale visual grounding datasets. In contrast, we focus on investigating the feasibility of effectively learning visual grounding under data-scarce settings.
Training with Synthetic Data. Building off image generation and image captioning, training on synthetic data has emerged as a promising research direction [51, 32, 87, 57, 80, 67, 28, 23, 70, 13]. In language tasks, research examines the generation of additional textual training data [17, 25, 48, 49, 72, 74, 89, 75, 56] or visual data conditioned on the text [65, 42, 19, 98, 44, 66, 85, 43] for data augmentation. In computer vision, [3, 91, 69] have demonstrated the benefits of augmenting real images with synthetic images as training data for image classification. In fine-grained understanding tasks, researchers also explore the use of synthetic data in object detection [15, 73, 7, 95, 16, 64], image segmentation [83, 38, 78, 14, 96], and visual grounding [22, 71]. [38, 95, 16] introduce a task-specific head into the image generator, leveraging intermediate-layer features of the generator. Once trained with the under the supervision of an off-the-shelf teacher model, the generator can synthesize images with annotation. In addition, some works [96, 14, 22] first generate images and then employ an auxiliary model to annotate the synthetic data. For instance, DiverGen [14] generate images by a vanilla diffusion model and subsequently depends on SAM [31] for dense annotation. Other methods [15, 7, 83, 73] employ visual priors (e.g., segmentation masks or edges) to guide the synthetic image generation process via controllable diffusion models. For example, FreeMask [83] uses FreestyleNet [81] to synthesize training images conditioned on target dataset-provided semantic masks. However, these methods rely on either the data generator, the annotator, or both to be pretrained on large-scale densely annotated datasets. In contrast, we leverages generative models trained solely on image-caption pairs, which make our method more flexible.
Data Selection. Recent advancements in data selection methods have gained significant attention from the research community [79, 27, 45]. These methods typically employ pre-defined criteria to assign a scalar score to each training example. While simple heuristics like repetition [33], string existence [53], or semantic similarity [1] are commonly used, more sophisticated criteria have also been explored, including gradient norms [52], forgetfulness [68], and influence function scores [84]. A common strategy is to prioritize data samples with lower training loss for subsequent training [88, 59, 8]. This approach leverages the observation that DNNs initially learn generalizable patterns before gradually overfitting to noisy data [2, 61, 37]. To counter confirmation bias, techniques such as co-training models [46, 20, 94, 34] have been proposed, where two networks are trained concurrently and use their combined predictions for selection. MentorNet [26] further extends this idea by introducing a learned curriculum to guide the selection of training data for a student network. In contrast to prior work, we introduce a novel filtering scheme tailored to our data-scarce setting. It select the most effective synthetic data from the hardness and overfitting perspectives.
6 Conclusions
In this paper, we introduce the POBF framework to effectively train a visual grounding model in challenging data-scarce scenarios. By leveraging advanced generative models pretrained solely on image-caption pairs, we synthesize training data efficiently. Specifically, we address the label misalignment issue by inpainting outside the bounding box. Additionally, to identify the most effective synthetic training samples, we propose a filtering scheme comprising three components: a hardness score, an overfitting score, and a penalty term. Experimental results show that our framework achieves an average improvement of 5.83% across four datasets, outperforming leading baselines with an accuracy increase ranging from 2.74% to 4.35%. Ablation studies highlight that our inpainting strategy and filtering scheme provide substantial accuracy gains of 3.53% and 1.96%, respectively. We further validate the robustness and generalizability of POBF across various generative models, data ratios, and model architectures. Finally, our framework sets a new direction for visual grounding in data-scarce settings, offering a scalable and efficient solution for real-world applications where annotated data is limited.